Solve the inequality | 4 - 2x | > 3. Give your solution in interval notation.
Answer by jim_thompson5910(35256)

| 4 - 2x | > 3


4 - 2x > 3 or 4 - 2x < -3


-2x > 3-4 or -2x < -3-4


-2x > -1 or -2x < -7


x < -1/(-2) or x > -7/(-2)


x < 1/2 or x > 7/2


=======================================================


The solution in interval notation is \left(-\infty,\frac{1}{2}\right)\cup\left(\frac{7}{2},\infty\right)